OpenStudy (anonymous):

Factor completely 169x2 - 64.

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @Squickers

TheSmartOne (thesmartone):

Do you know this formula? \(\sf (a+b)(a-b) = a^2-b^2\)

OpenStudy (anonymous):

oh yeah I learned that last year

TheSmartOne (thesmartone):

So we can also write \(\sf 169 = 13^2 \\ 64=8^2\)

OpenStudy (anonymous):

oh okay thanks that was easy (13x + 8)(13x - 8)

TheSmartOne (thesmartone):

Correct! :D

TheSmartOne (thesmartone):

Good job! :D

OpenStudy (anonymous):

thank you so much for your help

TheSmartOne (thesmartone):

Haha, anytime :) If you need help next time, you can tag me by putting that @ sign and then my username. So it would be like @TheSmartOne And I'll come and help you :)